A gas is stored at standard temperature and pressure (STP). The temperature is increased to 100° C while the volume remains unchanged. What is the new pressure of the gas?
1
Expert's answer
2020-12-24T06:32:23-0500

"\frac{P_1}{T_1} = \frac{P_2}{T_2}"


P2 = "\frac{P_1 × T_2}{T_1}"

P2 = P1 × "(\frac{T+100}{T}" )


so ratio "\frac{P_2}{P_1} = \frac{T+100}{T}"


whatever be the value of temperature accordingly pressure value will change.